# Modeling temporal hypergraphs

### v2

@@ 4 Model variants and other data
Most RHEM statistics discussed in this article have been defined via the prior degree $deg(I,E_{<t})$ counting the number of previous hyperedges containing $I$ as a subset. By this definition, the popularity of nodes – or the collective popularity of node sets – is independent of when, or how long ago, these past events happened. To allow that popularity may also fade out, we can introduce a temporal decay counting recent past events more strongly than prior events that happened in the distant past. A common way to define this is via an exponential decay in which the recent prior degree is defined by
@@
for a given half life period $T_{1/2}>0$ [65, 64, 66, 39]. An alternative to exponential decay is a sharp dropout of past events, once the time difference exceeds a given threshold [66]. Instead of fixing the shape and speed of a decay function, one may also seek to estimate it from the data [67, 68].
@@ 5 Conclusion
Appendix A Appendix: more detailed analysis of the Les Misérables data
In this paper, we illustrate the use of RHEM as a family of increasingly complex null models, applied to the Les Misérables data [52]. We estimate model parameters by maximum likelihood estimation with case-control sampling, where we bound the number of selected non-events per event to $\binom{80}{3}=82,160$. This means that for events $(t_{m},I_{m})$ of size one, two, and three we take the entire set of subsets $\binom{\mathcal{I}}{|I_{m}|}$ – compare the index set in the denominator of (3) – and for larger events we bound the set of sampled non-events to the number of different triads of nodes. Sampling non-events and computing hyperedge statistics is done with the eventnet software [39]. Given the table of hyperedge statistics, parameter estimation is done with the coxph function in the R package survival [42, 43].
Actors (nodes) in the Les Misérables data appear on average in $9.1$ chapters but their degree distribution is very skewed; compare the ten most frequent actors listed in Table 2. The sizes of hyperedges (i. e., the number of actors per chapter) varies from one to ten, where chapters with more than four actors are rare; see the histogram in Fig. 4.
Table 2: The ten most frequent actors in Les Misérables.
JV MA CO EN CR TH JA GA GI MY
#chapters: 113 77 55 31 27 27 24 21 20 19
Table 3: Parameters of six different RHEM estimated from the Les Misérables data. All models have been estimated with 288 events (chapters) and a cumulative number of $8,782,257$ selected non-events.
Model 1 Model 2 Model 3
[…]
subrep(3) $0.218\;(0.024)^{***}$
AIC $3935.734$ $3626.535$ $3493.831$
Model 4 Model 5 Model 6
[…]
subrep(3) $0.151\;(0.034)^{***}$
[…]
AIC $3558.687$ $3518.122$ $3489.317$
$^{***}p<0.001$; $^{**}p<0.01$; $^{*}p<0.05$
As discussed in the main text, temporal hypergraph models can give rise to a skewed activity distribution over the nodes by a preferential attachment mechanism, letting nodes with a larger number of past events be selected as participants of the next events with a higher probability. In general, preferential attachment for sets of order $k\geq 1$ is added to RHEM via the temporal hyperedge statistic subset repetition of order $k$, denoted by $subrep^{(k)}(I,E_{<t})$ and defined in Eq. (5).
We can use any estimated RHEM to compute the expected values of various quantities. For instance, according to Model 1 in Table 3, controlling only for preferential attachment of individual nodes, the three actors with the highest expected number of occurrences are actually the three most frequent actors in the observed data: JV (expected to appear in 121 chapters vs. 113 in the observed data), MA (expected 33; observed 77), and CO (expected 24; observed 55). We may explain the relative under-expectation of MA by the fact that MA appears for the first time in Volume 3, while JV, in contrast, appears throughout the novel at a more steady pace (compare Fig. 5). None of the models that we consider in this paper has a component that can explain why an actor does not appear at all for a long period and then suddenly becomes a rather frequent actor. In the main text we discussed the use of Model 1 and 2 in computing expected values for the number of joint appearances of actor pairs and triples.
Figure 5: Counting, for each chapter index $m=1,\dots,288$, the number of chapters with index $1,\dots,m$ featuring JV, MA, and CO, respectively.
Table 4: Parameters of RHEM including exact repetition estimated from the Les Misérables data. All models have been estimated with 288 events (chapters) and a cumulative number of $8,782,257$ selected non-events.
Model 3 Model 3 rep
[…]
subrep(3) $0.198\;(0.023)^{***}$ $0.196\;(0.023)^{***}$
repetition $0.075\;(0.018)^{***}$
AIC $3500.044$ $3483.827$
$^{***}p<0.001$; $^{**}p<0.01$; $^{*}p<0.05$
Table 5: Parameters RHEM including gender effects estimated from the Les Misérables data. All models have been estimated with 288 events (chapters) and a cumulative number of $8,782,257$ selected non-events.
Model 8 Model 9
female 0.184(0.097)-0.184\;(0.097) 0.126(0.114)-0.126\;(0.114)
homophily $0.212\;(0.046)^{***}$ $0.173\;(0.069)^{*}$
subrep(1) $0.054\;(0.003)^{***}$
subrep(2) $0.117\;(0.014)^{***}$
subrep(3) $0.134\;(0.035)^{***}$
closure 0.005(0.003)0.005\;(0.003)
assort(1) $0.011\;(0.002)^{***}$
assort(2) $0.005\;(0.002)^{**}$
AIC $4634.277$ $3411.816$
$^{***}p<0.001$; $^{**}p<0.01$; $^{*}p<0.05$
We can test a gender-specific activity effect via the statistic $female(I)$, defined in Eq. (9), and we can test homophily with respect to gender with the statistic $homophily(I)$, defined in Eq. (10). In the Les Misérables data, augmented by a gendergender variable as described in the main text, there are 28 female among the 80 actors. Thus, if actors were selected into chapters disregarding their gender (and not taking into account any other effect), we would expect that the average rate of females in chapters is $28/80$, that is 35%. In Table 5 we can see that female are not significantly overrepresented or underrepresented compared to this baseline ratio, but that there is homophily with respect to the “female” variable. This means that actors typically co-appear with other actors of the same gender, more than what would be expected by chance alone.

### v3

@@ 4 Model variants and other data
Most RHEM statistics discussed in this article have been defined via the prior degree $deg(I,E_{<t})$ counting the number of previous hyperevents containing $I$ as a subset. By this definition, the popularity of nodes – or the collective popularity of node sets – is independent of when, or how long ago, these past events happened. To allow that popularity may also fade out, we can introduce a temporal decay counting recent past events more strongly than prior events that happened in the distant past. A common way to define this is via an exponential decay in which the recent prior degree is defined by
@@
for a given half life period $T_{1/2}>0$ [67, 66, 68, 41]. The so-defined recent degree can then be used instead of the non-decaying degree in definitions of RHEM statistics. An alternative to exponential decay is a sharp dropout of past events, once the time difference exceeds a given threshold [68]. Instead of fixing the shape and speed of a decay function, one may also seek to estimate it from the data [69, 70].